OpenStudy (astrophysics):

Hey, I'm having some trouble with sketching solids with double integrals.

OpenStudy (astrophysics):

\[\int\limits \int\limits_R \sqrt{9-y^2} dA\] R = [0,4] x [0, 2]

OpenStudy (astrophysics):

I'm not really sure how to sketch this solid

OpenStudy (astrophysics):

 letting z = 9-y^2

OpenStudy (astrophysics):

\[z = \sqrt{9-y^2}\]

OpenStudy (anonymous):

notice that the defining function of the solid is independent of "x" and dependent on "y". and infact, a parabolic dependence on "y"

OpenStudy (anonymous):

let us re-write it as \[\int_{x=0}^4\int_{y=0}^2\sqrt{9-y^2}\,(dx\,dy)\]

OpenStudy (astrophysics):

Right

OpenStudy (astrophysics):

dydx no? :P

OpenStudy (anonymous):

moreover, the variables "x" and "y" are independent

OpenStudy (anonymous):

and "z" goes from \(-\infty\) to \(\infty\)

OpenStudy (anonymous):